OpenStudy (anonymous):

36degrees is the same as _____ radians. Round to the nearest hundredth of a radian.

OpenStudy (anonymous):

radians = \[36*\frac{ \pi }{ 180}\]

OpenStudy (anonymous):

so the answer is

OpenStudy (anonymous):

just do the math.\[\frac{ 36\pi }{ 180}\] Simplify then.

OpenStudy (anonymous):

.62

OpenStudy (anonymous):

.63 rounded, but yes.

OpenStudy (anonymous):

Good luck :)